OpenStudy (anonymous):

Solve the inequality 3(x + 6) 4x - 2 and show all of your work. You can use >= to represent the greater than or equal to symbol.

OpenStudy (anonymous):

Excuse me. what is the inequality sign of the question?

OpenStudy (anonymous):

>=

OpenStudy (anonymous):

3(x+6)>=4x-2 3x+18>=4x-2 20>=x x<=20